# ArkTS-CodeSearch: A Open-Source ArkTS Dataset for Code Retrieval

### v1

@@ Abstract.
ArkTS is a core programming language in the OpenHarmony ecosystem, yet research on ArkTS code intelligence is hindered by the lack of public datasets and evaluation benchmarks. This paper presents a large-scale ArkTS dataset constructed from open-source repositories, targeting code retrieval and code evaluation tasks. We design a single-search task, where natural language comments are used to retrieve corresponding ArkTS functions. ArkTS repositories are crawled from GitHub and Gitee, and comment–function pairs are extracted using tree-sitter-arkts, followed by cross-platform deduplication and statistical analysis of ArkTS function types. We further evaluate all existing open-source code embedding models on the single-search task and perform fine-tuning using both ArkTS and TypeScript training datasets, resulting in a high-performing model for ArkTS code understanding. This work establishes the first systematic benchmark for ArkTS code retrieval. Both the dataset and our fine-tuned model will be released publicly and are available at https://huggingface.co/hreyulog/embedinggemma_arkts and https://huggingface.co/datasets/hreyulog/arkts-code-docstring , establishing the first systematic benchmark for ArkTS code retrieval.
ArkTS, code embedding, benchmark, RAG
@@ 1. Introduction
In recent years, significant progress has been made in code semantic search, driven by large-scale datasets and pre-trained code embedding models. Benchmarks such as CodeSearchNet (Husain et al., 2019) have enabled systematic evaluation of natural language–to–code retrieval for mainstream programming languages including Python, Java, and JavaScript. These advances highlight the importance of high-quality datasets and well-defined evaluation protocols. Unfortunately, similar resources are not available for ArkTS, making it difficult to directly apply or fairly evaluate existing code retrieval methods in the OpenHarmony ecosystem.
To address this gap, we construct the first large-scale ArkTS dataset for code retrieval. Our dataset is collected from both GitHub [^1] https://github.com/ and Gitee [^2] https://gitee.com/, covering a wide range of OpenHarmony-related repositories. Using tree-sitter-arkts [^3] https://pypi.org/project/tree-sitter-arkts-open/, we parse ArkTS source code into abstract syntax trees and extract function- and class-level code units together with their associated natural language docstrings. The resulting dataset is carefully processed and reformatted into HuggingFace-compatible structures, facilitating reproducibility and future research.
Based on this dataset, we define a single-search code retrieval task, where natural language docstrings are used to retrieve corresponding ArkTS functions. We adopt the CodeSearchNet framework as a baseline retrieval protocol and conduct a comprehensive evaluation of existing open-source code embedding models on the ArkTS code search task. Furthermore, to better capture ArkTS-specific semantics, we fine-tune a pre-trained embedding model using ArkTS training data, achieving improved retrieval performance.
@@ 2.2. AST Parser
Abstract Syntax Tree (AST) parsers provide a structured and hierarchical representation of source code, explicitly encoding syntactic structure while implicitly capturing semantic relationships. AST-based representations have been widely adopted in code intelligence research, supporting tasks such as semantic code search, program analysis, and code summarization (Allamanis et al., 2018).
Compared with heuristic or regular-expression-based parsing, AST parsers offer improved robustness and syntactic correctness across diverse coding styles. Modern incremental parsers, such as Tree-sitter, enable scalable and language-aware code analysis and have been successfully applied to multiple programming languages. In this work, we leverage tree-sitter-arkts to parse ArkTS source files into ASTs, enabling accurate extraction of function- and class-level code units along with their associated docstrings. This structured preprocessing facilitates reliable code unit alignment and serves as a clean foundation for downstream code representation learning and evaluation.
@@ 2.3. Code Semantic Search
Code semantic search aims to retrieve relevant code snippets in response to a natural language query. Early approaches relied primarily on keyword matching or lexical similarity, which often fail to capture deeper semantic relationships between code and natural language. Recent advances leverage neural models to learn joint embedding spaces for code and text, significantly improving retrieval performance.
A range of pretrained models, such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), and CodeT5 (Wang et al., 2021), have demonstrated strong performance on semantic code search tasks by exploiting large-scale multilingual and multi-language code corpora. Benchmarks such as CodeSearchNet have further standardized evaluation protocols and accelerated progress in this area. However, these models and benchmarks predominantly focus on widely used languages, leaving emerging languages such as ArkTS underexplored. Consequently, there is a lack of systematic evaluation of code embedding models on ArkTS, highlighting the need for dedicated datasets and benchmarks to assess semantic code search performance in this setting.
Figure 1. Workflow for creating ArkTS-CodeSearch.
@@ 3.1.3. Code Parsing and Extraction
To extract structured code units, we employ tree-sitter-arkts to parse ArkTS source files into abstract syntax trees (ASTs). Based on the AST representation, we extract function- and class-level code units along with their associated docstrings. Each extracted record consists of a natural language docstring, the corresponding ArkTS function or class implementation, and its AST-based representation.
@@
3.2. CodeSearchNet
We adopt the CodeSearchNet framework as the baseline setting for ArkTS semantic code retrieval. In this paradigm, code search is formulated as a docstring-to-function retrieval task, where natural language docstrings serve as queries and function implementations are treated as retrieval targets. Both docstrings and code snippets are encoded into a shared embedding space, and relevant functions are retrieved based on vector similarity, typically using cosine similarity.
Following this formulation, each data instance in our ArkTS dataset consists of a function-level code snippet paired with its corresponding docstring.
As illustrated in Figure 2, we follow the CodeSearchNet paradigm by adopting a shared encoder for both docstrings and function code. The model is trained using contrastive learning to align semantically corresponding docstring–function pairs in a unified embedding space, and retrieval is performed via cosine similarity.
@@ 3.3. Evaluation Metrics
To assess retrieval performance, we adopt widely used metrics in code search literature:
@@ 3.3.1. MRR
Mean Reciprocal Rank (MRR) measures the average of the reciprocal ranks of the first relevant code snippet for each query (Craswell, 2009). It evaluates how quickly a model retrieves a correct function and is defined as:
@@
where $\text{rank}_{i}$ denotes the rank of the first relevant result for query $i$, and $|Q|$ is the total number of queries.
@@ 3.3.2. NDCG
Normalized Discounted Cumulative Gain (NDCG) (Manning et al., 2008) accounts for the position of relevant code snippets in the ranked list and is defined as:
@@
where $rel_{i}$ is the relevance of the result at position $i$, and $IDCG@k$ is the ideal DCG for normalization.
@@ 3.3.3. Recall
Recall@k measures the fraction of queries for which at least one relevant code snippet appears in the top $k$ retrieved results:
$\text{Recall@k}=\frac{\text{\# queries with relevant code in top k}}{|Q|}.$
3.4. Fine-Tune
@@
To enhance the semantic alignment between natural language descriptions and ArkTS code functions, we fine-tune a pre-trained code embedding model using a contrastive learning objective. Specifically, we adopt models to the ArkTS code retrieval task.
3.4.1. Training Data Construction.
Our training data is derived from the constructed ArkTS dataset, where each record contains a pair of a natural language docstring and its corresponding ArkTS function implementation. For each valid record, we generate paired training samples by treating the docstring as the query and the function body as the positive target. To further strengthen bidirectional semantic alignment, we also include the reverse pairs, where the function body is used as the query and the docstring as the target. Records with missing or empty fields are filtered out.
3.4.2. Model Architecture and Input Processing.
@@
We utilize the SentenceTransformer framework to load the pre-trained embedding model. Both docstrings and function bodies are encoded into dense vector representations using the same encoder. The maximum input sequence length is set to 512 tokens, and the output embedding dimension follows the original model configuration.
3.4.3. Training Objective.
The model is fine-tuned using a Multiple Negative Ranking Loss. Within each mini-batch, positive pairs are contrasted against other samples in the batch, which are implicitly treated as negative examples. This loss function encourages the model to assign higher similarity scores to true docstring–function pairs while pushing apart unrelated pairs. To improve efficiency on large batches, we employ a cached variant of the loss function.
3.4.4. Optimization Details.
We train the model for two epochs with a batch size of 32 and a learning rate of $2\times 10^{-5}$. The Adam optimizer is used with a linear warm-up schedule, where 10% of the total training steps are reserved for warm-up. Training samples are shuffled at each epoch to improve generalization.
3.4.5. Output Model.
After fine-tuning, the resulting embedding model is saved and used for downstream ArkTS code retrieval tasks. The fine-tuned model demonstrates improved semantic matching capability between natural language queries and ArkTS functions, serving as a strong baseline for ArkTS code intelligence applications.
3.4.6. Multiple Negative Ranking Loss
During fine-tuning, we employ the Multiple Negatives Ranking Loss, a contrastive learning objective widely used in dense retrieval and sentence embedding models (Reimers and Gurevych, 2019). This loss function encourages the model to discriminate a positive code snippet from multiple negative examples for the same query by leveraging in-batch negatives, without requiring explicit negative sampling.
@@ 3.4.6. Multiple Negative Ranking Loss
where $\mathrm{sim}(\cdot)$ denotes the cosine similarity between embeddings.
This objective can be interpreted as an instance of in-batch contrastive learning (van den Oord et al., 2019), which pulls semantically aligned docstring–function pairs closer in the embedding space while pushing apart mismatched pairs. Such a training strategy has been shown to be highly effective for semantic retrieval tasks, leading to more discriminative representations and improved code retrieval performance.
@@ 4. Dataset Statistics
We provide a comprehensive statistical analysis of our ArkTS dataset to better understand its characteristics. The dataset was collected on January 29, 2026, and is publicly available on HuggingFace [^5] https://huggingface.co/datasets/hreyulog/arkts-code-docstring.
@@
5. Experiments / Results
In this section, we evaluate the effectiveness of different embedding models on the ArkTS code retrieval task. We report results on the held-out test set using standard information retrieval metrics, including Mean Reciprocal Rank (MRR), NDCG@5, and Recall@K. All models are evaluated under the same retrieval setting to ensure fair comparison. The evaluation is conducted using our open-source retrieval evaluation framework, which is publicly available at https://github.com/hreyulog/retrieval_eval.
@@ 5.1. Leaderboard for opensorce models
Models Params MRR NDCG@5 Recall@1 Recall@5
Qwen/Qwen3-Embedding-0.6B 596M 0.6776 0.7015 0.6141 0.7723
google/embeddinggemma-300m 308M 0.6399 0.6654 0.5740 0.7416
BAAI/bge-m3 567M 0.5283 0.5603 0.4464 0.6558
BAAI/bge-base-zh-v1.5 110M 0.3598 0.3903 0.2841 0.4816
BAAI/bge-base-en-v1.5 110M 0.3439 0.3637 0.2935 0.4227
intfloat/e5-base-v2 110M 0.3073 0.3261 0.2596 0.3823
Table 1. Leaderboard for opensorce models
Table 1 reports the performance of several widely used open-source embedding models evaluated on the ArkTS code retrieval task without any task-specific fine-tuning.
@@
Among all evaluated models, Qwen/Qwen3-Embedding-0.6B (Zhang et al., 2025) achieves the best overall performance, with an MRR of 0.6776 and Recall@5 of 0.7723.
This indicates that large-scale general-purpose embedding models pretrained on multilingual and code-related corpora are more effective at capturing the semantic relationship between ArkTS docstrings and code.
@@
5.2. Fine-tuning on ArkTS Dataset
To improve retrieval performance, we fine-tune each embedding model on the ArkTS training set using a supervised contrastive learning objective. The results after fine-tuning are reported in Table 2. The best-performing fine-tuned model is released publicly on Hugging Face [^6] https://huggingface.co/hreyulog/embedinggemma_arkts to facilitate reproducibility and future research.
Overall, all evaluated models benefit substantially from fine-tuning, achieving consistent improvements across all evaluation metrics compared to their zero-shot counterparts. These gains confirm the effectiveness of task-specific supervision for aligning natural language docstrings with corresponding ArkTS code.
After fine-tuning, google/embeddinggemma-300m achieves the best overall performance, ranking first across all metrics with an MRR of 0.7788 and a Recall@5 of 0.8769. Notably, this model surpasses larger models such as Qwen/Qwen3-Embedding-0.6B, which ranks second with an MRR of 0.7745 and Recall@5 of 0.8643. This change in ranking indicates that mid-sized embedding models can achieve competitive or even superior performance when effectively adapted to the target domain.

### v2

@@ Abstract.
ArkTS is a core programming language in the OpenHarmony ecosystem, yet research on ArkTS code intelligence is hindered by the lack of public datasets and evaluation benchmarks. This paper presents a large-scale ArkTS dataset constructed from open-source repositories, targeting code retrieval and code evaluation tasks. We design a single-search task, where natural language comments are used to retrieve corresponding ArkTS functions. ArkTS repositories are crawled from GitHub and Gitee, and comment–function pairs are extracted using tree-sitter-arkts, followed by cross-platform deduplication and statistical analysis of ArkTS function types. We further evaluate existing open-source code embedding models on the single-search task and perform fine-tuning using both ArkTS and TypeScript training datasets, resulting in a high-performing model for ArkTS code understanding. This work establishes the first systematic benchmark for ArkTS code retrieval. Both the dataset and our fine-tuned model are available at https://huggingface.co/hreyulog/embedinggemma_arkts and https://huggingface.co/datasets/hreyulog/arkts-code-docstring.
ArkTS, code embedding, benchmark, RAG, Fine-tuning
@@ 1. Introduction
In recent years, significant progress has been made in code semantic search, driven by large-scale datasets and pre-trained code embedding models. Benchmarks such as CodeSearchNet (Husain et al., 2019) have enabled systematic evaluation of natural language–to–code retrieval for mainstream programming languages including Python, Java, and JavaScript, among others, along with subsequent datasets such as CodeXGLUE (Lu et al., 2021) and AdvTest, which further expand task coverage and evaluation settings.
To address this gap, we construct the first large-scale ArkTS dataset for code retrieval. Our dataset is collected from both GitHub [^1] https://github.com/ and Gitee [^2] https://gitee.com/, covering a wide range of OpenHarmony-related repositories. Using tree-sitter-arkts [^3] https://pypi.org/project/tree-sitter-arkts-open/, we parse ArkTS source code into abstract syntax trees (AST) and extract function-level code units together with their associated natural language docstrings. The resulting dataset is processed and reformatted into HuggingFace-compatible structures, facilitating reproducibility and future research.
Based on this dataset, we define a single-search code retrieval task, where natural language docstrings are used to retrieve corresponding ArkTS functions. We adopt the CodeSearchNet framework as a baseline retrieval protocol and conduct a comprehensive evaluation of existing open-source code embedding models on the ArkTS code search task. Furthermore, to better capture ArkTS-specific semantics, we fine-tune pre-trained embedding models using ArkTS training data, achieving improved retrieval performance.
@@ 2.2. AST Parser
AST parsers provide a structured and hierarchical representation of source code, explicitly encoding syntactic structure while implicitly capturing semantic relationships. AST-based representations have been widely adopted in code intelligence research, supporting tasks such as semantic code search, program analysis, and code summarization (Allamanis et al., 2018).
Compared with heuristic or regular-expression-based parsing, AST parsers offer improved robustness and syntactic correctness across diverse coding styles. Modern incremental parsers, such as Tree-sitter (Brunsfeld and others, 2020), enable scalable and language-aware code analysis and have been successfully applied to multiple programming languages. In this work, we leverage tree-sitter-arkts to parse ArkTS source files into ASTs, enabling accurate extraction of function- and class-level code units along with their associated docstrings. This structured preprocessing facilitates reliable code unit alignment and serves as a clean foundation for downstream code representation learning and evaluation.
@@ 2.3. Code Semantic Search
Code semantic search aims to retrieve relevant code snippets in response to a natural language query. Early approaches relied primarily on keyword matching or lexical similarity, which often fail to capture deeper semantic relationships between code and natural language. A commonly used method in this category is BM25(Robertson and Zaragoza, 2009), a sparse retrieval model based on term frequency–inverse document frequency (TF-IDF) scoring.
Recent advances leverage neural models to learn joint embedding spaces for code and text, significantly improving retrieval performance. A range of pretrained models, such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), and CodeT5 (Wang et al., 2021), have demonstrated strong performance on semantic code search tasks by exploiting large-scale multilingual and multi-language code corpora. Benchmarks such as CodeSearchNet have further standardized evaluation protocols and accelerated progress in this area. However, these models and benchmarks predominantly focus on widely used languages, leaving emerging languages such as ArkTS underexplored. Consequently, there is a lack of systematic evaluation of code embedding models on ArkTS, highlighting the need for dedicated datasets and benchmarks to assess semantic code search performance in this setting.
@@ 3.1.3. Code Parsing and Extraction
To extract structured code units, we employ tree-sitter-arkts to parse ArkTS source files into ASTs. Based on the AST representation, we extract function- and class-level code units along with their associated docstrings. Each extracted record consists of a natural language docstring, the corresponding ArkTS function or class implementation, and its AST-based representation.
@@
3.1.4. Licensing Considerations
The dataset was derived from open-source repositories hosted on GitHub and Gitee, which collectively encompass diverse license types including Apache, MIT, BSD, MPL, and others prevalent in the OpenHarmony ecosystem. The derived dataset – comprising function-docstring pairs extracted via AST parsing – constitutes a transformative derivative work intended for research purposes under fair use principles. However, users are advised to verify compliance with source repository licenses for any production or commercial deployment, as certain licenses may impose additional obligations (e.g., copyleft provisions).
3.2. Retrieval Methods
We evaluate two approaches for code retrieval. First, as a classic lexical retrieval method, BM25 ranks documents based on term frequency and inverse document frequency. For Chinese-heavy ArkTS docstrings, we segment both queries and code snippets using the jieba tokenizer  [^5] https://github.com/fxsjy/jieba, which combines dictionary-based matching with statistical modeling, constructs a directed acyclic graph of possible word sequences, and uses a Hidden Markov Model to handle new words. Despite its simplicity, BM25 provides a strong sparse baseline for comparison.
Second, semantic search leverages dense vector representations of queries and code snippets to capture semantic relationships beyond lexical overlap. Given a query $q$ and a code snippet $c$ , both are encoded into a shared embedding space using a pretrained code model, and their similarity is measured by cosine similarity (Reimers and Gurevych, 2019b):
$\text{sim}(q,c)=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{c}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{c}\|}.$
Code snippets are ranked according to similarity scores, enabling the retrieval of semantically relevant results even without direct lexical overlap, which is particularly important for ArkTS with bilingual or multilingual documentation.
We adopt the CodeSearchNet paradigm for ArkTS semantic code retrieval, formulating the task as docstring-to-function retrieval. Each data instance consists of a function-level code snippet paired with its corresponding docstring. As shown in Figure 2, a shared encoder is used for both docstrings and code, and the model is trained with contrastive learning to align semantically corresponding pairs in the embedding space. Retrieval is then performed via cosine similarity.
@@ 3.3. Evaluation Metrics
To assess retrieval performance, we adopt widely used metrics for retrieval-based code search. In our experiments, each query is searched against the entire test set to evaluate retrieval effectiveness.
@@ 3.3.1. MRR
Mean Reciprocal Rank (MRR) measures the average reciprocal rank of the first relevant code snippet for each query (Craswell, 2009). It reflects how early the correct function is retrieved and is defined as:
@@
where $\text{rank}_{i}$ denotes the rank position of the first relevant result for query $i$, and $|Q|$ is the total number of queries.
@@ 3.3.2. NDCG
Normalized Discounted Cumulative Gain (NDCG) (Manning et al., 2008) accounts for the positions of relevant code snippets in the ranked list:
@@
where $rel_{i}$ denotes the relevance score of the result at position $i$, and $IDCG@k$ is the ideal DCG for normalization.
@@ 3.3.3. Recall
Recall@k measures the proportion of queries for which the relevant code snippet appears within the top $k$ retrieved results:
$\text{Recall@k}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\mathbb{I}(\text{rank}_{i}\leq k),$
where $\mathbb{I}(\cdot)$ is the indicator function.
3.4. Fine-Tuning
@@
To enhance the semantic alignment between natural language descriptions and code functions, we fine-tune pre-trained code embedding models using a supervised contrastive learning objective, with the goal of improving retrieval performance on the ArkTS code search task.
3.4.1. Training Data Construction
For fine-tuning, we primarily use our ArkTS dataset, which provides function–docstring pairs suitable for supervised contrastive learning. In addition, we incorporate a large-scale TypeScript corpus introduced by Lozhkov et al. (2024), from which we extract 340,116 function–docstring pairs using tree-sitter-typescript parsing. All extracted pairs are processed using the same bidirectional alignment strategy as in ArkTS to ensure consistency.
We choose TypeScript as an auxiliary training source because ArkTS is a language derived from TypeScript, sharing similar syntax, type systems, and programming abstractions. As a result, TypeScript code provides semantically and structurally related supervision that can serve as an effective intermediate domain for model adaptation.
Based on these data sources, we explore three training configurations:
(1) ARKTS-TRAINING, which fine-tunes the model using only ArkTS pairs;
(2) TS-TRAINING, which fine-tunes the model using only the TypeScript-derived pairs; and
[…]
These configurations allow us to systematically study the impact of related-language pre-adaptation and domain-specific fine-tuning on ArkTS code retrieval performance.
3.4.2. Model Architecture and Input Processing
@@
We adopt the SentenceTransformer framework (Reimers and Gurevych, 2019b) to load and fine-tune pre-trained code embedding models. Docstrings and function bodies are encoded into dense vector representations using a shared encoder, ensuring that both modalities are mapped into the same embedding space. The maximum input length is set to 512 tokens, and the embedding dimensionality follows the configuration of the underlying pre-trained model.
3.4.3. Training Objective
Fine-tuning is conducted using the Multiple Negatives Ranking Loss (Reimers and Gurevych, 2019a), a contrastive learning objective widely adopted in dense retrieval. Within each batch, positive docstring–function pairs are contrasted against other examples in the batch, which are treated as implicit negatives. This encourages the model to assign higher similarity scores to true pairs while distancing unrelated ones. Conceptually, this can be viewed as a form of in-batch contrastive learning (van den Oord et al., 2019), which brings semantically aligned docstring–function pairs closer in the embedding space and pushes mismatched pairs further apart. Such a training strategy has been shown to be highly effective in semantic retrieval tasks, resulting in more discriminative representations and improved code retrieval performance.
@@ 3.4.3. Training Objective
where $\mathrm{sim}(\cdot)$ denotes cosine similarity. This in-batch contrastive objective encourages semantically aligned pairs to cluster in the embedding space while separating mismatched pairs (van den Oord et al., 2019).
3.4.4. Optimization Details
All models are trained for two epochs with a batch size of 32 and a learning rate of $2\times 10^{-5}$, using the Adam optimizer with a linear warm-up schedule for 10% of total steps. Training samples are shuffled each epoch to improve generalization.
@@ 4. Dataset Statistics
We provide a comprehensive statistical analysis of our ArkTS dataset to better understand its characteristics. The dataset was collected on January 29, 2026, and is publicly available on HuggingFace [^6] Available at https://huggingface.co/datasets/hreyulog/arkts-code-docstring.
@@
5. Experiments and Results
In this section, we evaluate the effectiveness of different embedding models on the ArkTS code retrieval task. We report results on the held-out test set using standard information retrieval metrics, including Mean Reciprocal Rank (MRR), NDCG@5, and Recall@K. All models are evaluated under the same retrieval setting to ensure fair comparison. The evaluation is conducted using our open-source retrieval evaluation framework, [^7] Available at https://github.com/hreyulog/retrieval_eval.
Models Params Original Finetuned (ARKTS-TRAINING)
MRR NDCG@5 Recall@1 Recall@5 MRR NDCG Recall@1 Recall@5
google/embeddinggemma-300m 308M 0.6399 0.6654 0.5740 0.7416 0.7788 0.8034 0.7142 0.8769
QWEN/QWEN3-EMBEDDING-0.6B 596M 0.6776 0.7015 0.6141 0.7723 0.7745 0.7971 0.7146 0.8643
BAAI/bge-m3 567M 0.5283 0.5603 0.4464 0.6558 0.7467 0.7729 0.6750 0.8504
BAAI/bge-base-zh-v1.5 110M 0.3598 0.3903 0.2841 0.4816 0.6748 0.7055 0.5961 0.7972
BAAI/bge-base-en-v1.5 110M 0.3439 0.3637 0.2935 0.4227 0.4866 0.5144 0.4195 0.5977
intfloat/e5-base-v2 110M 0.3073 0.3261 0.2596 0.3823 0.4855 0.5135 0.4158 0.5973
BM25 (jieba tokenizer) - 0.2043 0.2204 0.1643 0.2690 - - - -
Table 1. Comparison of original open-source models and finetuned models (ARKTS-TRAINING)
Models Params TS-TRAINING TS-ARKTS-TRAINING
MRR NDCG Recall@1 Recall@5 MRR NDCG Recall@1 Recall@5
google/embeddinggemma-300m 308M 0.6259 0.6525 0.5572 0.7318 0.8013 0.8229 0.7412 0.8867
Qwen/Qwen3-Embedding-0.6B 596M 0.6522 0.6785 0.5809 0.7563 0.7708 0.7947 0.7065 0.8655
BAAI/bge-m3 567M 0.6051 0.6338 0.5311 0.7191 0.7581 0.7846 0.6872 0.8630
BAAI/bge-base-zh-v1.5 110M 0.5270 0.5571 0.4509 0.6468 0.6759 0.7032 0.6018 0.7841
intfloat/e5-base-v2 110M 0.3751 0.3955 0.3226 0.4567 0.4557 0.4821 0.3904 0.5613
BAAI/bge-base-en-v1.5 110M 0.3834 0.4047 0.3299 0.4681 0.4539 0.4770 0.3949 0.5457
Table 2. Leaderboard for finetuned models under different training datasets (TS-TRAINING vs TS-ARKTS-TRAINING)
@@ 5.1. Leaderboard for Opensorce Models
Table 1 reports the performance of several widely used open-source embedding models evaluated on the ArkTS code retrieval task without any task-specific fine-tuning. In addition, we include a traditional lexical retrieval baseline, BM25 with the jieba tokenizer, for comparison.
@@
Among all evaluated models, Qwen/Qwen3-Embedding-0.6B (Zhang et al., 2025) achieves the best overall performance, with an MRR of 0.6776 and Recall@5 of 0.7723. This indicates that large-scale general-purpose embedding models pretrained on multilingual and code-related corpora are more effective at capturing the semantic relationship between ArkTS docstrings and code.
@@
We also observe that the BM25 baseline with jieba tokenization performs substantially worse than all embedding-based methods, achieving an MRR of 0.2043, Recall@1 of 0.2204, Recall@5 of 0.1643, and Recall@10 of 0.2690. This result suggests that purely lexical matching is insufficient for ArkTS code retrieval, where semantic alignment between docstrings and code snippets is crucial.
@@
5.2. Fine-Tuning